# The Infrared Absorption Spectrum of Phenylacetylene and its Deuterated Isotopologue in the Mid- to Far-IR

### 24 Jan 2024 v1

@@ IV Conclusions
Overall, very good agreement is present between experiment and theory for both phenylacetylene and d–phenylacetylene with an average difference of 1.2 and 1.0%, respectively, for the directly assigned features. The low-frequency rocking and bending motions at approximately 500 cm and below have some of the largest differences with an average difference of 2.5 and 2.7%. These modes are notoriously difficult to characterize with computational methods, and all of them deviate in a similar manner. Even so, an assignment is still easy to make for all features due to the sparsity of features in this region. As described in Section III.1, and 2 in phenylacetylene show the largest deviation between theory and experiment (3.8 and 6.1%) with the inaccuracy compounded for the overtone transition. This same deviation is not observed in d–phenylacetylene. The average difference on the order of 1% for these molecules is larger than that found for unsubstituted[18] and methyl-substituted[19] PAHs. This is not so surprising given the strong resonances present in phenylacetylene [16] and the known difficulty in computing vibrational spectra of acetylene molecules. [41]
Large isotopic frequency shifts are reported here for various modes that involve the acetylene hydrogen. Many of these modes either lose or gain intensity based on deuteration, causing major changes to the resultant absorption spectrum, particularly in the 500-900 cm (20-11.1 µm) range. Further experiments on isotopically labeled PAHs are needed, and such experiments are currently underway in this laboratory. Even with the insights provided herein, there remain quite a number of unanswered questions regarding how deuteration impacts the emission of PAHs, especially with regards to the strong infrared bands observed in the 1-20 µm range.[2, 1, 46] Due to a lack of experimental emission data for this wavelength region, further experimental and computational investigation is required to understand how anharmonicity and the isotopic frequency and intensity shifts impact the emission spectrum of standard and functionalized PAHs, molecules that are strongly tied to infrared observations pertinent to astronomical observation during the age of JWST and its possible successors.
@@ Acknowledgements.
- [2] C. Boersma, Louis J. Allamandola, Vincent J. Esposito, Alexandros Maragkoudakis, J. D. Bregman, P. Temi, Timothy J. Lee, Ryan C. Fortenberry, and Els Peeters. JWST: Deuterated PAHs, PAH-nitriles, and PAH Overtone and Combination Bands I: Program Description and First Look. ApJ, In Press, 2023.
@@
- [30] Mark H Stockett, James N Bull, Henrik Cederquist, Suvasthika Indrajith, MingChao Ji, José E Navarro Navarrete, Henning T Schmidt, Henning Zettergren, and Boxing Zhu. Efficient stabilization of cyanonaphthalene by fast radiative cooling and implications for the resilience of small pahs in interstellar clouds. Nature Communications, 14(1):395, 2023.

### 22 Feb 2024 v2

@@ IV Conclusions
Overall, very good agreement is present between experiment and theory for both phenylacetylene and d–phenylacetylene with an average difference in frequency of 1.2 and 1.0%, respectively, for the directly assigned features. The low-frequency rocking and bending motions at approximately 500 cm and below have some of the largest differences in frequency with an average difference of 2.5 and 2.7%. These modes are notoriously difficult to characterize with computational methods, and all of them deviate in a similar manner. Even so, an assignment is still easy to make for all features due to the sparsity of features in this region. As described in Section III.1, and 2 in phenylacetylene show the largest deviation in frequency between theory and experiment (3.8 and 6.1%) with the inaccuracy compounded for the overtone transition. This same deviation is not observed in d–phenylacetylene. The average difference on the order of 1% for these molecules is larger than that found for unsubstituted[18] and methyl-substituted[19] PAHs. This is not so surprising given the strong resonances present in phenylacetylene [16] and the known difficulty in computing vibrational spectra of acetylene molecules. [43]
Large isotopic frequency shifts are reported here for various modes that involve the acetylene hydrogen. Many of these modes either lose or gain intensity based on deuteration, causing major changes to the resultant absorption spectrum, particularly in the 500-900 cm (20-11.1 µm) range. Further experiments on isotopically labeled PAHs are needed, and such absorption experiments are currently underway in our laboratory. Even with the insights provided herein, there remain quite a number of unanswered questions regarding how deuteration impacts the emission of PAHs, especially with regards to the strong infrared bands observed in the 1-20 µm range.[2, 1, 49] Very little PAH, and their deuterated isotopologues in particular, emission data exist in the literature. A recent emission experiment on phenylacetylene and d–phenylacetylene carried out by Lacinbala et al.,[50] and a follow-up computational study by our group[16] look into the role that anharmonicity plays in the emission of substituted aromatic molecules. Outside of this recent example, there exists a dearth of experimental emission data for this wavelength region, requiring further experimental and computational investigation to understand how anharmonicity and the isotopic frequency and intensity shifts impact the emission spectrum of standard and functionalized PAHs, molecules that are strongly tied to infrared observations pertinent to astronomical observation during the age of JWST and its possible successors.
V Supplementary Material
The data that support the findings of this study are available in the supplemental information and from the corresponding author upon reasonable request.
@@ Acknowledgements.
- [2] C Boersma, L J Allamandola, V J Esposito, A Maragkoudakis, J D Bregman, P Temi, T J Lee, R C Fortenberry, and E Peeters. JWST: Deuterated PAHs, PAH Nitriles, and PAH Overtone and Combination Bands. I. Program Description and First Look. The Astrophysical Journal, 959(74), December 2023.
@@
- [32] Vasyl Yatsyna, Daniël J. Bakker, Peter Salén, Raimund Feifel, Anouk M. Rijs, and Vitali Zhaunerchyk. Infrared Action Spectroscopy of Low-Temperature Neutral Gas-Phase Molecules of Arbitrary Structure. Physical Review Letters, 117(11):118101, September 2016.
- [33] Sander Jaeqx, Jos Oomens, Alvaro Cimas, Marie-Pierre Gaigeot, and Anouk M. Rijs. Gas-Phase Peptide Structures Unraveled by Far-IR Spectroscopy: Combining IR-UV Ion-Dip Experiments with Born–Oppenheimer Molecular Dynamics Simulations. Angewandte Chemie International Edition, 53(14):3663–3666, 2014. _eprint: https://onlinelibrary.wiley.com/doi/pdf/10.1002/anie.201311189.
@@
- [39] Vincenzo Barone. Anharmonic vibrational properties by a fully automated second-order perturbative approach. The Journal of Chemical Physics, 122(1):014108, January 2005.
@@
- [48] Ryan C. Fortenberry, Timothy J. Lee, and Holger S. P. Müller. Excited vibrational level rotational constants for SiC2: A sensitive molecular diagnostic for astrophysical conditions. Molecular Astrophysics, 1:13–19, November 2015.
@@
- [50] Ozan Lacinbala, Géraldine Féraud, Julien Vincent, and Thomas Pino. Aromatic and Acetylenic C–H or C–D Stretching Bands Anharmonicity Detection of Phenylacetylene by UV Laser-Induced Vibrational Emission. The Journal of Physical Chemistry A, 126(30):4891–4901, August 2022.